### v1

# Class-preserving Coleman Automorphisms of Finite Groups with Semidihedral Sylow 2-Subgroups

Abstract.
In this paper, we prove that finite groups with semidihedral Sylow 2-subgroup have Class-preserving Coleman outer automorphism group of odd order. As a consequence, these groups satisfy the normalizer problem. In particular, we extend some existing results in the literature concerning class-preserving Coleman automorphisms of finite groups with semidihedral Sylow 2-subgroups.
Key words and phrases:
Class-preserving automorphism, Coleman automorphism, Semidihedral 2-group, Sylow 2-subgroup, normalizer problem
2010 Mathematics Subject Classification:
20D45, 20D20, 16U70
The author’s ORCiD: $0$-0001-8834-4358
The author is member of INdAM-GNSAGA (Italy).
The author declares that he has no conflict of interest.
1. Introduction
The study of class-preserving automorphisms of finite groups – that is, automorphisms that preserve the conjugacy classes of the elements of the group – has long been a subject of interest in group theory research. Since the early twentieth century, several researchers have investigated which finite groups admit only inner class-preserving automorphisms, or have sought to construct counterexamples to this property. For relevant references, see for example [5, 11, 12, 20, 24].
In this paper we focus on the study of class-preserving automorphisms that satisfy an additional property: they become inner when restricted to every Sylow $p$-subgroup. Automorphisms satisfying this property are called Coleman automorphisms, due to their significance in the study of the normalizer of a finte group $G$ in the unit group $\mathcal{U}(\mathbb{Z}(G))$ of its integral group ring $\mathbb{Z}G$, as shown by Coleman in [7, Theorem 1].
More precisely, let $G$ be a finite group. We denote by $\operatorname{Aut}_{c}(G)$ the subgroup of $\operatorname{Aut}(G)$ consisting of all class-preserving automorphisms of $G$; by $\operatorname{Aut}_{Col}(G)$, the subgroup of $\operatorname{Aut}(G)$ consisting of all Coleman automorphisms of $G$; and by $\operatorname{Aut}_{\mathbb{Z}}(G)$, the subgroup of $\operatorname{Aut}(G)$ consisting of all automorphisms of $G$ that induce an inner automorphism on the integral group ring $\mathbb{Z}G$. From Coleman’s result [7, Theorem 1], together with a result of Jackowski and Marciniak [18, Proposition 2.3], it follows that $\operatorname{Aut}_{\mathbb{Z}}(G)\leq\operatorname{Aut}_{c}(G)\cap\operatorname{Aut}_{Col}(G)$. Moreover, Krempa proved that $\operatorname{Aut}_{\mathbb{Z}}(G)$ is an elementary abelian 2-group. Therefore, denoting by $\operatorname{Out}_{c}(G)=\operatorname{Aut}_{c}(G)/\operatorname{Inn}(G)$ the class-preserving outer automorphism group and by $\operatorname{Out}_{Col}(G)=\operatorname{Aut}_{Col}(G)/\operatorname{Inn}(G)$ the Coleman outer automorphism group, we deduce that if $G$ is a finite group such that
$\operatorname{Out}_{c}(G)\cap\operatorname{Out}_{Col}(G)\text{ is of odd order,}$, (1)
then $\operatorname{Aut}_{\mathbb{Z}}(G)=\operatorname{Inn}(G)$.
It is worth noting that classifying finite groups $G$ for which $\operatorname{Aut}_{\mathbb{Z}}(G)=\operatorname{Inn}(G)$ is equivalent to solving the well known normalizer problem [21, Problem 43], that is,
$N_{\mathcal{U}(\mathbb{Z}G)}(G)=GZ(\mathcal{U}(\mathbb{Z}G)).$
In the literature, one can find several affirmative results and counterexamples concerning the problem of identifying finite groups $G$ that satisfy Property (1) (see, e.g., [13, 16, 19, 23, 25]).
The aim of this paper is to prove that any finite group with semidihedral Sylow 2-subgroups satisfies Property (1), and consequently satisfies the normalizer problem. In particular, we extend [25, Theorem 3.5], where the authors prove that for any solvable finite group $G$ with semidihedral Sylow 2-subgroups, $\operatorname{Out}_{c}(G)\cap\operatorname{Out}_{Col}(G)$ has odd order. Moreover, considering a different definition of Coleman automorphisms given in [19] – namely, using our notation, those automorphisms $\varphi\in\operatorname{Aut}_{c}\cap\operatorname{Aut}_{Col}(G)$ such that $\varphi^{2}\in\operatorname{Inn}(G)$ – it follows that any such automorphism of odd order is inner. Hence, the main theorem of this paper also extends the results proved in [2].
2. Preliminary results
In this section, we state the main theorem of the paper and present some preliminary results from [12, 14, 16] that are useful for its proof.
Let $G$ be a finite group. Denote by $\operatorname{Aut}(G)$ the group of all automorphisms of $G$, and by $\operatorname{Inn}(G)$ the subgroup of $\operatorname{Aut}(G)$ consisting of all inner automorphisms. Define the subgroup $\operatorname{Aut}_{c}(G)\leq\operatorname{Aut}(G)$ as the set of all class-preserving automorphisms of $G$, i.e., those automorphisms that map each element of $G$ to a conjugate of itself. An automorphism $\varphi\in\operatorname{Aut}(G)$ is called a Coleman automorphism if, for every Sylow $p$-subgroup $S$ of $G$, we have $\varphi|_{S}=\operatorname{\mathrm{conj}}_{g}|_{S}$ for some $g\in G$, where $\operatorname{\mathrm{conj}}_{g}$ denotes the inner automorphism defined by $\operatorname{\mathrm{conj}}_{g}(x)=g^{-1}xg$. We denote by $\operatorname{Aut}_{Col}(G)$ the group of Coleman automorphisms of $G$. Notice that both $\operatorname{Aut}_{c}(G)$ and $\operatorname{Aut}_{Col}(G)$ contain $\operatorname{Inn}(G)$. Accordingly, we define $\operatorname{Out}_{c}(G)=\operatorname{Aut}_{c}(G)/\operatorname{Inn}(G)$ and $\operatorname{Out}_{Col}(G)=\operatorname{Aut}_{Col}(G)/\operatorname{Inn}(G)$ as the class-preserving outer automorphism group and the Coleman outer automorphism group.
The aim of this paper is to prove the following result.
Theorem 2.1.
If $G$ is a finite group with semidihedral Sylow 2-subgroup, then $G$ satisfies Property (1); in other words, $\operatorname{Out}_{c}(G)\cap\operatorname{Out}_{Col}(G)$ is of odd order.
In order to prove the main theorem, we will need the following results. As usual, we denote by $O_{p}(G)$ the maximal normal $p$-subgroup of $G$, which is the intersection of all Sylow $p$-subgroups of $G$, and by $O_{p^{\prime}}(G)$ the maximal normal subgroup of $G$ whose order is coprime to $p$.
Lemma 2.2.
Let $N\unlhd G$ and let $p$ be a prime number which does not divide the order of $G/N$. Then the following hold:
  • (i)
    If $\varphi\in\operatorname{Aut}(G)$ is a class-preserving or a Coleman automorphism of $G$ of $p$-power order, then $\varphi$ induces a class-preserving or a Coleman automorphism of $N$, respectively.
  • (ii)
    If $\operatorname{Out}_{c}(N)$ or $\operatorname{Out}_{Col}(N)$ is a $p^{\prime}$-group, then so is $\operatorname{Out}_{c}(G)$ or $\operatorname{Out}_{Col}(G)$, respectively. Moreover, if $\operatorname{Out}_{c}(N)\cap\operatorname{Out}_{Col}(N)$ is a $p^{\prime}$-group, then so is $\operatorname{Out}_{c}(G)\cap\operatorname{Out}_{Col}(G)$.
Proof.
See [12, Corollary 5] and [16, Corollary 3]. ∎
Proposition 2.3.
Let $G$ be a semidirect product $K\rtimes H$, where $K$ is an elementary abelian $p$-group, for some prime $p$. Let $\varphi\in\operatorname{Aut}(G)$ be defined by $h\varphi=h$ for all $h\in H$, and $k\varphi=k^{m}$ for all $k\in K$ and for some fixed positive integer $m$. If $\varphi\in\operatorname{Aut}_{c}(G)$, then $\varphi$ induces a class-preserving automorphism of $N_{G}(U)/U$, for each subgroup $U$ of index $p$ in $K$.
Proof.
Since $\operatorname{Aut}_{c}(G)=\operatorname{Aut}_{\mathbb{Q}}(G)$ (see [18, Proposition 2.8]), setting $R=\mathbb{Q}$, the claim holds from [14, Proposition 2.8]. ∎
Lemma 2.4.
[12, Lemma 5] Let $p$ be a prime, and let $G$ be a finite group such that there is a non-inner automorphism of $G$ of $p$-power order which preserves the conjugacy classes of elements of $p$-power order, but induces inner automorphisms on all proper factor groups. Then the following hold.
  • (i)
    The Frattini subgroup $\Phi(G)$ is a $p$-group.
  • (ii)
    If Sylow $p$-subgroups of $G$ are abelian, then $O_{p}(G)=1$.
  • (iii)
    If $O_{p}(G)=1$, and $C_{G}(N)\leq N$ for some $1\neq N\unlhd G$, then $\varphi|_{N}\neq\gamma|_{N}$ for all $\gamma\in\operatorname{Inn}(G)$.
Finally, we will need the following well-known lemma, whose proof can be found in [12, Lemma 2].
Lemma 2.5.
Let $p$ be a prime, and an automorphism of $G$ of $p$-power order. Assume further that there is $N\unlhd G$ such that $\varphi$ fixes all elements of $N$,and that $\varphi$ induces the identity on $G/N$. Then $\varphi$ induces the identity on $G/O_{p}(Z(N))$. If $\varphi$ fixes in addition a Sylow $p$-subgroup of $G$ element-wise, then $\varphi$ is an inner automorphism.
Moreover, throughout the paper, the following remark will play an important role.
Remark 1.
[12, Remark 1] Let $\varphi$ be a non-inner (class-preserving) automorphism of $G$, of order a power of $p$. Assume that
  • (i)
    there is $U\leq G$ and $\sigma\in\operatorname{Inn}(G)$ with $\varphi|_{U}=\sigma|_{U}$, or that
  • (ii)
    there is $N\unlhd G$ with $N\varphi=N$ such that $\varphi$ induces an inner automorphism on $G/N$.
Then there is $\gamma\in Inn(G)$ such that (i) $\varphi\gamma|_{U}=\mathrm{id}$ or (ii) $\varphi\gamma$ induces the identity on $G/N$, and $\varphi\gamma$ is still a non-inner (class-preserving) automorphism of order a power of $p$. Sometimes we will use this fact without any further comment, just saying that we “modify by an inner automorphism”, when we intend to replace $\varphi$ by $\varphi\gamma$.
From now on, we will consider $G$ a finite group with semidihedral Sylow 2-subgroups. Let us recall that a semidihedral group $S$ of order $2^{n}$ is a group with the following presentation
$SD_{2^{n}}=\langle a,b\mid a^{2^{n-1}}=b^{2}=1,bab=a^{2^{n-2}-1}\rangle,$
or, in other words, is the semidirect product $C_{2^{n-1}}\rtimes C_{2}$, where $C_{2}$ acts on $C_{2^{n-1}}$ by $x\mapsto x^{2^{n-2}-1}$. In particular, all the elements in $S\setminus\langle a\rangle$ are $a^{2i}b$, for $10\leq i\leq 2^{n-2}-$ 1, of order 2, or $a^{2i+1}b$, for $10\leq i\leq 2^{n-2}-$ 1, of order 4. Moreover, we have the following useful properties:
  • (i)
    $S$ possesses precisely three maximal subgroups, respectively the cyclic group $C=\left\langle a\right\rangle$, the generalized quaternion group $Q=\left\langle a^{2},ab\right\rangle$, and the dihedral group $D=\left\langle a^{2},b\right\rangle$. More generally, the subgroups of $S$ are cyclic, generalized quaternion, dihedral or Klein four-groups;
  • (ii)
    the center of $S$ is $Z(S)=\left\langle a^{2^{n-2}}\right\rangle$, which has order 2, and $S/Z(S)\cong D_{2^{n-2}}$, the dihedral group of order $2^{n-1}$;
  • (iii)
    the maximal cyclic subgroups of $C$, $D$ and $Q$ all concide with $\Phi(S)=[S,S]=<a^{2}>\cong C_{2^{n-2}}$, and we have $S/\Phi(S)=S/[S,S]\cong C_{2}\times C_{2}$ ;
  • (iv)
    the normal subgroups of $S$ are the three maximal normal subgroups $C$, $D$ and $Q$, togheter with all the subgroups of $C$. In particular, every proper normal subgroup of $S$ is either maximal or is cyclic and contained in $[S,S]$.
Notice that, by definition, $n\geq 4$, or, in other words, $S$ has order at least $16$. For more details on the properties of a semidihedral group, see, e.g., [1, Lemma 1, p. 9].
3. Proof of Theorem 2.1
In this section, we prove Theorem 2.1, beginning with several lemmas needed to show that any finite group with a semidihedral Sylow 2-subgroup satisfies Property (1).
Throughout this section, we assume that $G$ is a minimal counterexample to Theorem 2.1. In particular, we suppose that $G$ admits a class-preserving, non-inner Coleman automorphism of order a power of 2, which we denote by $\varphi$, without further mention. Since a Sylow 2-subgroup of $G$ is semidihedral, and its subgroups, together with the subgroup itself, are cyclic, dihedral, or generalized quaternion, then the Sylow 2-subgroups of subgroups and quotients of $G$ are of these types. Hence, by the minimality of $G$, [14, Proposition 4.7], and the main theorem of [15], proper quotient, and proper subquotient of $G$ satisfies Property $\mathcal{W}$.
Let us denote by $S$ a Sylow 2-subgroup of $G$, and let $a,y\in G$ be such that $\left\langle a\right\rangle$ is the maximal cyclic subgroup of $S$ not containing $y)$. In other words, $y$ has order 2 or 4.
Remark 2.
Notice that any proper quotient $G/N$ of $G$ must have even order. Otherwise, $N$ would contain a Sylow 2-subgroup $S=SD_{2^{n}}$, and by Lemma 2.2, since $\varphi$ is assumed to be a non-inner class-preserving Coleman automorphism of order a power of 2, $\varphi$ would induce a class-preserving Coleman automorphism on $N$ of order a power of 2. This contradicts the minimality of $G$. In particular, no proper normal subgroup of $G$ contains $S$.
Let $F=F(G)$ denote the Fitting subgroup of $G$, and let $F^{*}=F^{*}(G)=FE$ be the generalized Fitting subgroup of $G$, where $E=E(G)$ is the layer of $G$. See [9, Chapter 1, Section 4] for more details on the properties of $F$ and $F^{*}$ that we will use in this paper. For instance, $F=\prod_{p\in\pi(G)}O_{p}(G)$, $F^{*}$ is a self centralizing subgroup and is a central product of $F$ and $E$, and $Z(E)$ is an abelian normal subgroup of $G$ contained in $F$.
The proof of Theorem 2.1 relies on several auxiliary lemmas.
Lemma 3.1.
If $G$ is a minimal counterexample to Theorem 2.1, then $O_{2^{\prime}}(F)$ is non-trivial.
Proof.
Suppose, by way of contradiction, that $F$ is a 2-group; that is, $F=O_{2}(G)$. By Remark 2, we can modify $\varphi$ by an inner automorphism so that it fixes $S$ element-wise.
If $F=F^{*}$, then $C_{G}(F)\leq F$. Since $F$ is a 2-group, it follows that $F\leq S$ and so $F$ is fixed element-wise by $\varphi$. Therefore, $\varphi$ induces the identity on $G/F$, and hence, by Lemma 2.5, $\varphi\in\operatorname{Inn}(G)$, which is a contradiction.
Hence, we can assume $E\neq 1$. Suppose $Z(E)=1$, it then follows that $F^{*}=F\times E$. If $E$ contains a cyclic Sylow 2-subgroup, then by the Burnside’s normal $p$-complement theorem [6, Theorem II, Section 243], it has a normal 2-complement, and hence, by Feit-Thompson theorem [8], it is solvable. This is impossible since $E$ is a product of non-abelian simple groups. Therefore, $E$ cannot contain a cyclic Sylow 2-subgroup. Since a Sylow 2-subgroup of $F^{*}=F\times E$ is a subgroup of the semidihedral group $S$, and a Sylow 2-subgroup of $E$ cannot be cyclic, it follows that $O_{2}(G)=F=1$, and $F^{*}=E$ is a non-abelian simple group. In [3, 4, 22], Brauer and Suzuki proved that a generalized quaternion cannot be a Sylow 2-subgroup of a finite simple group. Moreover, by Remark 2, a Sylow 2-subgroup of $E$ cannot be a semidihedral group, since $E$ is a normal subgroup of $G$. Therefore a Sylow 2-subgroup of $E$ can be only a dihedral group. In [10] it was proved that the only finite simple groups with a dihedral Sylow 2-subgroup are $\mathrm{PSL}_{2}(q)$, and $\mathrm{A}_{7}$. Noting that every Coleman automorphism of a group is, up to an inner automorphism, a $p$-central automorphism for any prime $p$ dividing the order of the group, it follows from [16, Theorem 14] that there exists a prime $p\in\pi(E)$ such that all $p$-central automorphisms of $E$ are inner. So, also in this case, we can modify $\varphi$ so that $E=F^{*}$ is fixed element-wise. Consequently, $\varphi$ induces the identity on $G/F^{*}$. Since $Z(F^{*})=Z(E)=1$, it follows that $\varphi=\mathrm{id}$, which is a contradiction. Thus we have $1\neq Z(E)\leq F$ is a 2-group.
Let $P$ be a Sylow 2-subgroup of $E<F^{*}=FE$. As shown above, $P$ cannot be cyclic. Moreover, $P$ cannot be a Klein four-group, since it is a subgroup of a semidihedral group and $F$ is a 2-group, hence it is non-abelian. Consequently, $Z(E)=Z(P)\cong C_{2}$. On the one hand, a Sylow 2-subgroup $P/Z(P)$ of $F^{*}/Z(E)$ is a dihedral group, since $P$ is non-cyclic subgroup of a semidihedral group. On the other hand, since, by Remark 2, a Sylow 2-subgroup of $G$ cannot be a normal subgroup, we conclude that if $F\neq 1$, then $O_{2}(G)=F$ must be either a cyclic group or a dihedral group or a generalized quaternion group. Recalling that $F^{*}=FE$ such that $[F,E]=1$ and $F\cap E=Z(E)$, it follows that the Sylow 2-subgroup of $F^{*}/Z(E)$ can be a dihedral group only if $F=1$; that is, $F^{*}=E$. As above, we can modify $\varphi$ so that it fixes $F^{*}$ element-wise. Moreover, we can suppose that $\varphi$ stabilizes $S$ and that $P\leq S$. Hence, since $\varphi$ is a Coleman automorphism, there exists $s\in S$ such that $\varphi|_{S}=\operatorname{\mathrm{conj}}_{s}|_{S}$ and in particular $[P,s]=1$. If $s\not\in P$, then $\left\langle P,s\right\rangle$ is a non-abelian 2-group with center of order at least 4, which is a contraciction. Otherwise, if $s\in P$, then $s\in Z(P)=Z(E)=Z(F^{*})$. This implies that we can assume $\varphi|_{F^{*}}=\mathrm{id}$ and also $\varphi|_{S}=\mathrm{id}$, since $s\in Z(P)=Z(S)$ and $\varphi|_{S}=\operatorname{\mathrm{conj}}_{s}|_{S}$. By Lemma 2.5, it follows that $\varphi$ is an inner automorphism, which is a contradiction. Therefore, $O_{2^{\prime}}(F)\neq 1$. ∎
Lemma 3.2.
If $G$ is a minimal counterexample to Theorem 2.1, then the Frattini subgroup $\Phi(G)$ of $G$ is a 2-group. In particular, the subgroup $O_{2^{\prime}}(F)$ is a product of abelian minimal normal subgroup of $G$ and it has a complement in $G$.
Proof.
Note that $G$ satisfies the hipotheses of Lemma 2.4, since it is a counterexample of Theorem 2.1. Therefore $\Phi(G)$ is a 2-group. By well-known properties of the Frattini subgroup and the Fitting subgroup (see, e.g. [17, Chapter III, 3.14, 4.4, 4.5]), since $\Phi(G)$ is a 2-group, and so $O_{2^{\prime}}(G)\cap\Phi(G)=1$, the second part of the claim also follows. ∎
Lemma 3.3.
Let $G$ be a minimal counterexample to Theorem 2.1. If $M$ is a minimal normal subgroup of $G$ contained in $O_{2^{\prime}}(G)$, then $M$ has a complement $K$ in $G$, and it is possible to modify $\varphi$ so that $K$ is fixed element-wise by $\varphi$. Moreover, if $N\neq M$ is a normal subgroup of $G$, then there exists a 2-element $k\in K$ such that
  • $k$ inverts the elements of $M$;
  • the image of $k$ in $G/MN$ is a central involution; and
  • $\varphi|_{M}=\operatorname{\mathrm{conj}}_{k}|_{M}$.
Proof.
Since $M$ is a minimal normal subgroup contained in $O_{2^{\prime}}(G)$, it follows from Lemma 2.2 that $M\varphi=M$; moreover, by Lemma 3.2, $M$ has a complement $K$ in $G$, and by [12, Proposition 1], $K\varphi=K$. Since $K\cong G/M$, the minimality of $G$ implies that $\varphi$ induces an inner automorphism on $K$. Hence, by Remark 1, it is possible to modify $\varphi$ so that $K$ is fixed element-wise by $\varphi$. Moreover, if $N\neq M$ is a normal subgroup of $G$, then, again by the minimality of $G$, $\varphi$ induces an inner automorphism on $G/N$. Since $G=M\rtimes K$ and $M$ is an abelian minimal normal subgroup of $G)$, such an inner automorphism is a conjugation by a 2-element $k\in K$, since $\varphi$ has order a power of 2. In particular, $\varphi|_{M}=\operatorname{\mathrm{conj}}_{k}|_{M}$.
Note that, for each $g\in G$,
$g^{-1}kgMN=(kMN)\varphi=k\varphi MN=kMN,$
and so the image of $k$ is central in $G/MN$. Moreover, $\varphi^{2}|_{M}=\operatorname{\mathrm{conj}}_{k^{2}}|_{M}$ has order a power of 2 and $M\leq O_{2^{\prime}}(G)$, it follows that 2 does not divide the order of $M$, and so it does not divide the cardinality of any conjugacy class in $M$. Thus $M=\{m\in M\mid m\varphi^{2}=m\}$, and so $\varphi^{2}|_{M}=\mathrm{id}$. Furthermore, $C_{M}(k)=C_{M}(\varphi)\unlhd G$, and so, since $M$ is minimal normal subgroup of $G$, we must have either $C_{M}(\varphi)=M$ or $C_{M}(\varphi)=1$. Since $\varphi$ does not act trivially on $M$, we conclude that $C_{M}(\varphi)=1$. Finally, $k$ acts by inversion on $M$, since $[k^{2},M]=1$. ∎
Lemma 3.4.
Let $G$ be a minimal counterexample to Theorem 2.1, and let $M$ be a minimal normal subgroup of $G$ contained in $O_{2^{\prime}}(G)$ with complement $K$ in $G$. Suppose that $\varphi$ fixes $K$ element-wise. Assume that $M$ is not a cyclic group, and let $U$ be a maximal subgroup of $M$. Then $\varphi$ induces an inner automorphism on $N_{G}(U)/U$.
Proof.
Note that $M$ is a subgroup of a Sylow $p$-subgroup $Q$ of $G$. Since $\varphi$ is a Coleman automorphism, we have $\varphi|_{Q}=\operatorname{\mathrm{conj}}_{g}|_{Q}$ for some $g\in G$. As $M$ is normal in $G$, it follows that for each $m\in M\leq Q$, we have $m\varphi=g^{-1}mg\in M$. Moreover, as shown above, $g$ acts on $M$ by inversion. Therefore, $g$ normalizes $U$, and $\varphi$ induces a Coleman automorphism $\psi$ on $N_{G}(U)/U$. By Proposition 2.3, the automorphism $\psi$ is also class-preserving, and hence, by the minimality of $G$, it must be inner. ∎
Lemma 3.5.
Let $G$ be a minimal counterexample to Theorem 2.1. Then any proper cyclic quotient of $G$ has order 2.
Proof.
Let $N\unlhd G$ be such that $G/N$ is cyclic. By Remark 2, it follows that $N$ does not contain $S$. Then a Sylow 2-subgroup $Q$ of $G/N$ is cyclic and must be contained in the abelianization $S/[S,S]$ of $S$. Since $S/[S,S]\cong C_{2}\times C_{2}$, it follows that $Q\cong C_{2}$. If $G/N$ has order strictly greater than 2, then $\varphi$ induces a class-preserving Coleman automorphism on $G/N$ of order a power of 2, which is clearly non-inner, since $G/N$ is cyclic. This contradicts [14, Proposition 4.7]. ∎
Lemma 3.6.
Let $G$ be a minimal counterexample to Theorem 2.1, and let $E$ be the layer of $G$. Then $O_{2}(G)E$ is non-trivial.
Proof.
Suppose, by way of contradiction, that $O_{2}(G)E=1$. This implies that $F^{*}=F$ is a $2^{\prime}$-group and, by Lemma 3.2, it is a direct product of abelian minimal normal subgroups $M_{i}$ of $G$, with $1\leq i\leq\ell$. Note that $\ell>1$. Indeed, if $F=M_{1}$, since $C_{G}(F)=C_{G}(F^{*})\leq F^{*}=F$, by Lemma 2.4(iii), it follows that $\varphi|_{M_{1}}\neq\gamma|_{M_{1}}$ for each $\gamma\in\operatorname{Inn}(G)$, which contradicts Lemma 3.3.
Denote by $M_{\hat{i}}$ the subgroup of $G$ generated by $\{M_{j}\mid 1\leq j\leq\ell,\ j\neq i\}$. By Lemma 3.2 and Lemma 3.3, there exists a complement $H$ to $F$, and there are involutions $h_{i}\in H$ acting on $M_{i}$ by inversion. Moreover, we can modify $\varphi$, up to an inner automorphism, so that it fixes $M_{\hat{i}}H$ element-wise, and $\varphi|_{M_{i}}=\operatorname{\mathrm{conj}}_{h_{i}}|_{M_{i}}$.
Suppose $h_{i}\neq h_{j}$, for some $i\neq j$. Note that only one $h_{\bar{i}}$ can belong to $Z(S)$, and we can suppose $h_{i},h_{j}\not\in Z(S)$. Then $S=\left\langle yh_{j},h_{i}\right\rangle$, for some $y\in H\setminus\left\langle a\right\rangle$ of order 4. Thus, $FS$ is a normal subgroup of $G$, and, by Remark 2, $G=FS$. This contradicts [25, Theorem 3.5]. Thus, $h_{i}=h_{j}$ for all $1\leq i,j\leq\ell$.
Now we fix an index $i$, and assume that $M_{i}$ is cyclic, generated by $m_{i}$. Let $K=C_{H}(M_{i})$. Note that $M_{i}\unlhd H$ implies $C_{H}(M_{i})\unlhd H$. Moreover, since $M_{i}$ is a cyclic minimal normal subgroup of odd order, it follows that $M_{i}\cong C_{p}$ for some prime $p$. Hence, $H/K$ is isomorphic to a subgroup of $C_{p-1}$, and so it is cyclic. Furthermore, $h_{i}$ does not belong to $K$, since it acts by inversion on $M_{i}$, and so, by Lemma 3.5, $H/K\cong C_{2}$, generated by $h_{i}$. Thus, we conclude that $H=K\times\langle h_{i}\rangle$. Note that, since $\left\langle h_{i}\right\rangle$ is normal in $H$, and hence normal in $S$, and since $S$ is semidihedral, it follows that the order of $K$ is not divisible by 4. By Lemma 3.3, we can modify $\varphi$ so that it fixes $M_{i}H$ element-wise. Let $v\in M_{\hat{i}}$ and $k\in K$. There exists $g\in G$ – in particular, it is enough to take $g\in M_{\hat{i}}K\langle h_{i}\rangle$ – such that $(m_{i}vk)\varphi=g^{-1}(m_{i}vk)g$. This implies that $(vk)\varphi=g^{-1}(vk)g$ and $m_{i}=m_{i}\varphi=g^{-1}m_{i}g$, and so, by the latter equality, $g\in M_{\hat{i}}K$. Therefore, $\varphi|_{M_{\hat{i}}K}$ is a class-preserving automorphism of $M_{\hat{i}}K$; hence, by [16, Proposition 4.6], since 2 divides the order of $K$ but 4 does not, $\varphi|_{M_{\hat{i}}K}$ must be an inner automorphism. Thus, there exists $k\in K$ such that $\varphi|_{M_{\hat{i}}}=\operatorname{\mathrm{conj}}_{k}|_{M_{\hat{i}}}$, and so $\varphi|_{F}=\operatorname{\mathrm{conj}}_{k}|_{F}$, since $K$ commutes with the elements of $M_{i}$. This contradicts Lemma 2.4(iii), and therefore $M_{i}$ cannot be cyclic.
Now we modify $\varphi$ so that it fixes $M_{\hat{1}}K$ element-wise and $\varphi|_{M_{1}}=\operatorname{\mathrm{conj}}_{k_{1}}|_{M_{1}}$, and we choose a maximal subgroup $U$ of $M_{1}$. By Lemma 3.4, $\varphi$ induces an inner automorphism on $N_{G}(U)/U$. Then there exists a non-trivial 2-element $h\in N_{H}(U)$ such that $[h,M_{\hat{1}}]=1$ and $h$ acts on $M_{1}/U$ by inversion. Substituting $S$ with a suitable conjugate, if necessary, we may suppose $h\in S$. The same reasoning applied to another minimal normal subgroup shows that there exists a non-trivial 2-element $k\in H$ such that $[k,M_{1}]=1$, and we may suppose that $k\in S$. As above, $h_{1}$ is a central involution in $S$, and we may assume that it belongs to the maximal cyclic subgroup $C=\langle a\rangle$ of $S=SD_{2^{n}}$. Since $h_{1}$ acts on $M_{1}$ by inversion and $M_{1}$ has no elements of order 2, it follows that $h_{1}$ acts fixed-point free on $M_{1}$. In particular, $h_{1}=a^{2^{n-2}}$. Moreover, $a$ also acts fixed-point free on $M_{1}$, and so we may suppose that $k\in\left\langle a\right\rangle$, that is, $k=a^{s}y$, for some $s\in\mathbb{N}$. Also, $h=a^{t}y$, for some $t\in\mathbb{N}$, since $[h,M_{\hat{1}}]=1$ and $h_{1}$ – which equals $h_{i}$ for all $i$ – acts fixed-point free on $M_{\hat{1}}$. Then $hk^{-1}=a^{t-s}$. Since $h=a^{t-s}k$, and $k$ fixes $M_{1}$ element-wise, it follows that $h$ acts on $M_{1}$ by conjugation with $a^{t-s}$. Since $h$ acts on $M_{1}/U$ by inversion, it follows that $a^{t-s}=h_{1}$. Therefore, $h=a^{t-s}k=h_{1}k$, and this implies that $\varphi|_{F}=\operatorname{\mathrm{conj}}_{h}|_{F}$, which contradicts Lemma 2.4(iii). ∎
Lemma 3.7.
Let $G$ be an abelian minimal counterexample to Theorem 2.1. Then $O_{2}(G)$ is non-trivial.
Proof.
Suppose, by way of contradiction, that $O_{2}(G)=1$. This implies that the order of $F$ is odd, and by Lemma 3.6, $E\neq 1$. Moreover, $Z(E)$ also has odd order. However, since $Z(E)=\Phi(E)\leq\Phi(G)$, and $\Phi(G)$ is a 2-group, it follows that $Z(E)=1$. Therefore, $E$ is a non-abelian simple group. By Lemma 3.2, there exists an abelian minimal normal subgroup of $O_{2^{\prime}}(F)$ which has a complement $K$ in $G$. Recalling that $M\leq O_{2^{\prime}}(F)\leq F$ and $F\cap E=1=M\cap K$, we have $E\leq K$. Moreover, we can suppose that $\varphi$ fixes $K$ element-wise and acts on $M$ by inversion.
Suppose that $M$ is not cyclic. Let $U$ be a maximal subgroup of $M$. By Lemma 3.4, $\varphi$ induces an inner automorphism on $N_{G}(U)/U$. Hence, there exists a 2-element $g\in K$ such that $(h\varphi)h^{-g}\in U$ for each $h\in N_{G}(U)$. In particular, since $E\leq K$ and $\varphi$ fixes $K$ element-wise, we have $[E,g]=1$, and $g$ acts on $M/U$ by inversion. Let $P$ be a Sylow 2-subgroup of $E$. As shown in the proof of Lemma 3.1, $P$ is not cyclic, $[P,g]=1$, and $\langle P,g\rangle$ is a 2-group with center of order greater than or equal to 4, which is impossible. Therefore, $M$ is cyclic, and, by Lemma 3.5, $K/C_{K}(M)\cong C_{2}$.
By Lemma 3.3, there exists a 2-element $k\in K$ which maps to a central involution $\bar{k}$ in $\bar{K}=K/E$, and which acts on $M$ by inversion. Moreover, $K/E=\langle\bar{k}\rangle\times(C_{K}(M)/E)$.
Assume that $C_{K}(M)=E$, that is, $G/ME=\left\langle\bar{k}\right\rangle\cong C_{2}$. Let $1\neq m\in M$ and $e\in E$. Then there exists $c\in K$ such that $(me)\varphi=(me)^{c}$, that is, $m^{-1}=m\varphi=c^{-1}mc$ and $e=e\varphi=c^{-1}ec$. It follows that $c=hk$, for some $h\in E$, and so $\varphi|_{E}=\operatorname{\mathrm{conj}}_{k}|_{E}$ is a class-preserving automorphism of $E$. Hence, $\operatorname{\mathrm{conj}}_{k}|_{E}\in\operatorname{Inn}(E)$, considering the structure of $E$ described in the proof of Lemma 3.1 and recalling that a class-preserving automorphism, up to an inner automorphism, is $p$-central for every prime $p$ dividing the order of $E$. In other words, there exists a 2-element $e\in E$ such that $\operatorname{\mathrm{conj}}_{k}|_{E}=\operatorname{\mathrm{conj}}_{e}|_{E}$, and, in particular, $[k,e]=1$. This implies that $ke^{-1}$ is a 2-element not belonging to $E$, and it centralizes a Sylow 2-subgroup of $E$. Note that this leads to a contradiction: indeed, $E$ must be either a dihedral or a generalized quaternion group, and the subgroup generated by $ke^{-1}$ centralizes a Sylow 2-subgroup of $E$ while having trivial intersection with $E$. However, this is not possible, since the Sylow 2-subgroups of $G$ are semidihedral. Therefore, we conclude that $C_{K}(M)/E\neq 1$.
Let $P$ be a Sylow 2-subgroup of $C_{K}(M)/E$. Then $P$ is non-trivial, since, by Remark 2, $E$ does not contain $S$. Since $K\cong C_{2}\times(C_{K}(M)/E)$, again by the structure of a semidihedral group, it follows that $P$ has order 2. Hence, a Sylow 2-subgroup of $G/E=MK/E$ is an elementary abelian 2-group of order 4, and so it is $S/[S,S]$. It follows that a Sylow 2-subgroup of $E$ is $[S,S]$, which is cyclic. However, $E$ cannot contain a cyclic Sylow 2-subgroup, and so this leads to a contradiction. Therefore, we conclude that $O_{2}(G)=1$. ∎
Lemma 3.8.
Let $G$ be a minimal counterexample to Theorem 2.1, and let $S=SD_{2^{n}}$ be a semidihedral Sylow $2$-subgroup of $G$. Then $Z(S)$ is normal in $G$. In particular, $Z(S)=O_{2}(G)$
Proof.
First note that, by Lemma 3.7 and Remark 2, $O_{2}(G)$ is a non-trivial proper normal subgroup of $S$.
Suppose that $O_{2}(G)$ is the maximal dihedral subgroup $D$ of $S$ or the maximal generalized quaternion subgroup $Q$ of $S$. Note that $Z(S)=Z(D)=Z(Q)$, and since the center is characteristic in both $D$ and $Q$, it follows that $Z(S)$ is normal in $G$ in either case.
Now suppose that $O_{2}(G)$ is one of the cyclic subgroups contained in the maximal cyclic subgroup $C$ of $S$. If $O_{2}(G)=Z(S)$, we are done. Otherwise, in each of the other cases – recalling that every subgroup of a cyclic group is characteristic – it follows again that $Z(S)$ is normal in $G$.
Finally, suppose $O_{2}(G)\neq Z(S)$. It follows that $O_{2}(G)$ has order at least 4. Let $z\in O_{2}(G)\setminus Z(S)$ and let $1\neq m\in M$. Then there exists a 2-element $g\in K$ such that $(mz)\varphi=(mz)^{g}$, that is, $m^{-1}=m\varphi=m^{g}$ and $z=z\varphi=z^{g}$. It follows that $\left\langle O_{2}(G),g\right\rangle$ is a non-cyclic abelian 2-group of order at least 8, which is impossible. Hence, $O_{2}(G)=Z(S)$. ∎
Before stating the following lemma, we recall some previously established properties of a minimal counterexample to Theorem 2.1. By Lemma 3.1 and Lemma 3.3, there exists a minimal normal subgroup of $G$ contained in $O_{2^{\prime}}(F)$, which has a complement $K$ in $G$. Modifying $\varphi$ according to Lemma 3.3, we may assume that $\varphi$ fixes $K$ element-wise. Furthermore, by Lemma 3.7 and again Lemma 3.3, there exists a 2-element $\tilde{k}\in K$ – which may be taken to lie in $S\leq K$ – such that $[\tilde{k},G]\leq MO_{2}(G)$, and both $\varphi$ and $\tilde{k}$ act on $M$ by inversion. Note that, by Lemma 3.4 and Lemma 3.8, we also conclude that $[\tilde{k},G]\leq MZ(S)$; in particular, $[\tilde{k},S]\leq Z(S)$.
Lemma 3.9.
Let $G$ be a minimal counterexample to Theorem 2.1, and let $M$ be a minimal normal subgroup of $G$ contained in $O_{2^{\prime}}(F)$. Then $M$ is cyclic.
Proof.
The structure of the semidihedral group $S=SD_{2^{n}}$ implies that if $1\neq[k,S]\leq Z(S)$, then $\tilde{k}\in\left\langle a\right\rangle$ and has order 4; that is, $\tilde{k}=a^{2^{n-3}}$. Let $s=y$ be an element of order 2 or 4 in $S\setminus C$, where $C$ is the maximal normal cyclic subgroup of $S$, and let $P=\langle\tilde{k},s\rangle$. Note that $Z(P)=Z(S)$.
Assume, by way of contradiction, that $M$ is not cyclic. Note that $s^{2}=1$ or $s^{2}=a^{2^{n-2}}\in Z(S)$; hence $[s^{2},M]=1$, since, by Lemma 3.8, $Z(S)\cong C_{2}$ is normal in $G$. In particular, there exists a maximal subgroup $U$ of $M$ such that $P$ normalizes $U$, that is, $P\leq N_{K}(U)$. By Lemma 3.4, $\varphi$ induces an inner automorphism on $N_{K}(U)M/U$; thus, there exists a 2-element $g\in N_{K}(U)$ such that $(h\varphi)h^{-g}\in U$ for all $h\in N_{K}(U)M$. In particular, $[g,P]=1$, and so $\left\langle P,g\right\rangle$ is a 2-group. If $g\in P$, then $g\in Z(P)=Z(S)$, and so $g$ would act trivially on $M$; however, this is impossible since, by Lemma 3.4, $g$ acts on $M$ by inversion on $M/U$. If $g\notin P$, then $\left\langle P,g\right\rangle$ would be a non-abelian 2-group with center of order at least 4, which is a contradiction, since $Z(P)=Z(S)\cong C_{2}$. Therefore, we conclude that $M$ is cyclic. ∎
Now we are ready to conclude the proof of Theorem 2.1.
Proof of Theorem 2.1.
Let us consider the notations introduced above, and let $G$ be a minimal counterexample to Theorem 2.1. By Lemma 3.9, a minimal normal subgroup $M$ of $G$ contained in $O_{2^{\prime}}(F)$ is cyclic; in particular, $M\cong C_{p}$ for some prime $p$. It follows that $K/C_{K}(M)$ is cyclic, and thus, by Lemma 3.5, we have $K/C_{K}(M)\cong C_{2}$.
By the beginning of the proof of Lemma 3.9, we know that $\tilde{k}\in K$ belongs to the cyclic subgroup of order 4 in $S$, which is normal in $S$ and lies in $[S,S]\cong C_{2^{n-2}}$. This leads to a contradiction, since $K/C_{K}(M)\cong C_{2}$, and so its Sylow 2-subgroups must be isomorphic to either $S/C_{2^{n-1}}$, $S/D_{2^{n-2}}$, or $S/Q_{2^{n-1}}$. In all these cases, $[S,S]\leq C_{K}(M)$, implying that $K$ acts trivially on $M$. However, we know that $\tilde{k}\in K$ acts on $M$ by inversion. Hence, we obtain a contradiction.
This completes the proof of the main theorem. ∎
Acknowledgements
The author acknowledges the funding support from MUR-Italy via PRIN 2022RFAZCJ "Algebraic methods in Cryptanalysis".