### v2

# Class-preserving Coleman Automorphisms of Finite Groups with Semidihedral Sylow 2-Subgroups

Abstract.
In this paper, we prove that finite groups with semidihedral Sylow 2-subgroup have Class-preserving Coleman outer automorphism group of odd order. As a consequence, these groups satisfy the normalizer problem. In particular, we extend some existing results in the literature concerning class-preserving Coleman automorphisms of finite groups with semidihedral Sylow 2-subgroups.
Key words and phrases:
Class-preserving automorphism, Coleman automorphism, Semidihedral 2-group, Sylow 2-subgroup, normalizer problem
2010 Mathematics Subject Classification:
20D45, 20D20, 16U70
The author’s ORCiD: 0000-0001-8834-4358
The author is member of INdAM-GNSAGA (Italy).
The author declares that he has no conflict of interest.
1. Introduction
The study of class-preserving automorphisms of finite groups – that is, automorphisms that preserve the conjugacy classes of the elements of the group – has long been a subject of interest in group theory research. Since the early twentieth century, several researchers have investigated which finite groups admit only inner class-preserving automorphisms, or have sought to construct counterexamples to this property. For relevant references, see for example [5, 11, 12, 21, 25].
In this paper we focus on the study of class-preserving automorphisms that satisfy an additional property: they become inner when restricted to every Sylow $p$-subgroup. Automorphisms satisfying this property are called Coleman automorphisms, due to their significance in the study of the normalizer of a finte group $G$ in the unit group $\mathcal{U}(\mathbb{Z}(G))$ of its integral group ring $\mathbb{Z}G$, as shown by Coleman in [7, Theorem 1].
More precisely, let $G$ be a finite group. We denote by $\operatorname{Aut}_{c}(G)$ the subgroup of $\operatorname{Aut}(G)$ consisting of all class-preserving automorphisms of $G$; by $\operatorname{Aut}_{Col}(G)$, the subgroup of $\operatorname{Aut}(G)$ consisting of all Coleman automorphisms of $G$; and by $\operatorname{Aut}_{\mathbb{Z}}(G)$, the subgroup of $\operatorname{Aut}(G)$ consisting of all automorphisms of $G$ that induce an inner automorphism on the integral group ring $\mathbb{Z}G$. From Coleman’s result [7, Theorem 1], together with a result of Jackowski and Marciniak [18, Proposition 2.3], it follows that $\operatorname{Aut}_{\mathbb{Z}}(G)\leq\operatorname{Aut}_{c}(G)\cap\operatorname{Aut}_{Col}(G)$. Moreover, Krempa [18, Theorem 3.2] proved that $\operatorname{Aut}_{\mathbb{Z}}(G)$ is an elementary abelian 2-group. Therefore, denoting by $\operatorname{Out}_{c}(G)=\operatorname{Aut}_{c}(G)/\operatorname{Inn}(G)$ the class-preserving outer automorphism group and by $\operatorname{Out}_{Col}(G)=\operatorname{Aut}_{Col}(G)/\operatorname{Inn}(G)$ the Coleman outer automorphism group, we deduce that if $G$ is a finite group such that
$\operatorname{Out}_{c}(G)\cap\operatorname{Out}_{Col}(G)\text{ is of odd order,}$ (1)
then $\operatorname{Aut}_{\mathbb{Z}}(G)=\operatorname{Inn}(G)$.
It is worth noting that classifying finite groups $G$ for which $\operatorname{Aut}_{\mathbb{Z}}(G)=\operatorname{Inn}(G)$ is equivalent to solving the well known normalizer problem [22, Problem 43], that is,
$N_{\mathcal{U}(\mathbb{Z}G)}(G)=GZ(\mathcal{U}(\mathbb{Z}G)).$
For further details on the normalizer problem, see, for example, [19].
In the literature, one can find several affirmative results (see, e.g., [16, 20, 24, 26]) and a counterexample provided by Hertweck in [13] concerning the problem of identifying finite groups $G$ that satisfy Property (1).
The aim of this paper is to prove that any finite group with semidihedral Sylow 2-subgroups satisfies Property (1), and consequently satisfies the normalizer problem. In particular, we extend [26, Theorem 3.5], where the authors prove that for any solvable finite group $G$ with semidihedral Sylow 2-subgroups, $\operatorname{Out}_{c}(G)\cap\operatorname{Out}_{Col}(G)$ has odd order. Moreover, considering a different definition of Coleman automorphisms given in [20] – namely, using our notation, those automorphisms $\varphi\in\operatorname{Aut}_{c}\cap\operatorname{Aut}_{Col}(G)$ such that $\varphi^{2}\in\operatorname{Inn}(G)$ – it follows that any such automorphism of odd order is inner. Hence, the main theorem of this paper also extends the results proved in [2].
2. Preliminary results
In this section, we state the main theorem of the paper and present some preliminary results from [12, 14, 16] that are useful for its proof.
Let $G$ be a finite group. Denote by $\operatorname{Aut}(G)$ the group of all automorphisms of $G$, and by $\operatorname{Inn}(G)$ the subgroup of $\operatorname{Aut}(G)$ consisting of all inner automorphisms. Define the subgroup $\operatorname{Aut}_{c}(G)\leq\operatorname{Aut}(G)$ as the set of all class-preserving automorphisms of $G$, i.e., those automorphisms that map each element of $G$ to a conjugate of itself. An automorphism $\varphi\in\operatorname{Aut}(G)$ is called a Coleman automorphism if, for every Sylow $p$-subgroup $S$ of $G$, we have $\varphi|_{S}=\operatorname{\mathrm{conj}}_{g}|_{S}$ for some $g\in G$, where $\operatorname{\mathrm{conj}}_{g}$ denotes the inner automorphism defined by $\operatorname{\mathrm{conj}}_{g}(x)=g^{-1}xg$. We denote by $\operatorname{Aut}_{Col}(G)$ the group of Coleman automorphisms of $G$. Notice that both $\operatorname{Aut}_{c}(G)$ and $\operatorname{Aut}_{Col}(G)$ contain $\operatorname{Inn}(G)$. Accordingly, we define $\operatorname{Out}_{c}(G)=\operatorname{Aut}_{c}(G)/\operatorname{Inn}(G)$ and $\operatorname{Out}_{Col}(G)=\operatorname{Aut}_{Col}(G)/\operatorname{Inn}(G)$ as the class-preserving outer automorphism group and the Coleman outer automorphism group.
The aim of this paper is to prove the following result.
Theorem 2.1.
If $G$ is a finite group with semidihedral Sylow 2-subgroup, then $G$ satisfies Property (1); in other words, $\operatorname{Out}_{c}(G)\cap\operatorname{Out}_{Col}(G)$ is of odd order.
In order to prove the main theorem, we will need the following results. As usual, we denote by $O_{p}(G)$ the maximal normal $p$-subgroup of $G$, which is the intersection of all Sylow $p$-subgroups of $G$, and by $O_{p^{\prime}}(G)$ the maximal normal subgroup of $G$ whose order is coprime to $p$.
Lemma 2.2.
Let $N\unlhd G$ and let $p$ be a prime number which does not divide the order of $G/N$. Then the following hold:
  • (i)
    If $\varphi\in\operatorname{Aut}(G)$ is a class-preserving or a Coleman automorphism of $G$ of $p$-power order, then $\varphi$ induces a class-preserving or a Coleman automorphism of $N$, respectively.
  • (ii)
    If $\operatorname{Out}_{c}(N)$ or $\operatorname{Out}_{Col}(N)$ is a $p^{\prime}$-group, then so is $\operatorname{Out}_{c}(G)$ or $\operatorname{Out}_{Col}(G)$, respectively. Moreover, if $\operatorname{Out}_{c}(N)\cap\operatorname{Out}_{Col}(N)$ is a $p^{\prime}$-group, then so is $\operatorname{Out}_{c}(G)\cap\operatorname{Out}_{Col}(G)$.
Proof.
See [12, Corollary 5] and [16, Corollary 3]. ∎
Proposition 2.3.
Let $G$ be a semidirect product $K\rtimes H$, where $K$ is an elementary abelian $p$-group, for some prime $p$. Let $\varphi\in\operatorname{Aut}(G)$ be defined by $h\varphi=h$ for all $h\in H$, and $k\varphi=k^{m}$ for all $k\in K$ and for some fixed positive integer $m$. If $\varphi\in\operatorname{Aut}_{c}(G)$, then $\varphi$ induces a class-preserving automorphism of $N_{G}(U)/U$, for each subgroup $U$ of index $p$ in $K$.
Proof.
Since $\operatorname{Aut}_{c}(G)=\operatorname{Aut}_{\mathbb{Q}}(G)$ (see [18, Proposition 2.8]), setting $R=\mathbb{Q}$, the claim holds from [14, Proposition 2.8]. ∎
Lemma 2.4.
[12, Lemma 5] Let $p$ be a prime, and let $G$ be a finite group such that there is a non-inner automorphism of $G$ of $p$-power order which preserves the conjugacy classes of elements of $p$-power order, but induces inner automorphisms on all proper factor groups. Then the following hold.
  • (i)
    The Frattini subgroup $\Phi(G)$ is a $p$-group.
  • (ii)
    If Sylow $p$-subgroups of $G$ are abelian, then $O_{p}(G)=1$.
  • (iii)
    If $O_{p}(G)=1$, and $C_{G}(N)\leq N$ for some $1\neq N\unlhd G$, then $\varphi|_{N}\neq\gamma|_{N}$ for all $\gamma\in\operatorname{Inn}(G)$.
Finally, we will need the following well-known lemma, whose proof can be found in [12, Lemma 2].
Lemma 2.5.
Let $p$ be a prime, and an automorphism of $G$ of $p$-power order. Assume further that there is $N\unlhd G$ such that $\varphi$ fixes all elements of $N$,and that $\varphi$ induces the identity on $G/N$. Then $\varphi$ induces the identity on $G/O_{p}(Z(N))$. If $\varphi$ fixes in addition a Sylow $p$-subgroup of $G$ element-wise, then $\varphi$ is an inner automorphism.
Moreover, throughout the paper, the following remark will play an important role.
Remark 1.
[12, Remark 1] Let $\varphi$ be a non-inner (class-preserving) automorphism of $G$, of order a power of $p$. Assume that
  • (i)
    there is $U\leq G$ and $\sigma\in\operatorname{Inn}(G)$ with $\varphi|_{U}=\sigma|_{U}$, or that
  • (ii)
    there is $N\unlhd G$ with $N\varphi=N$ such that $\varphi$ induces an inner automorphism on $G/N$.
Then there is $\gamma\in Inn(G)$ such that (i) $\varphi\gamma|_{U}=\mathrm{id}$ or (ii) $\varphi\gamma$ induces the identity on $G/N$, and $\varphi\gamma$ is still a non-inner (class-preserving) automorphism of order a power of $p$. Sometimes we will use this fact without any further comment, just saying that we “modify by an inner automorphism”, when we intend to replace $\varphi$ by $\varphi\gamma$.
Notation 2.6.
From now on, we will consider $G$ a finite group with semidihedral Sylow 2-subgroups. Let us recall that a semidihedral group $S$ of order $2^{n}$ is a group with the following presentation
$SD_{2^{n}}=\langle a,b\mid a^{2^{n-1}}=b^{2}=1,bab=a^{2^{n-2}-1}\rangle,$
or, in other words, is the semidirect product $C_{2^{n-1}}\rtimes C_{2}$, where $C_{2}$ acts on $C_{2^{n-1}}$ by $x\mapsto x^{2^{n-2}-1}$. In particular, all the elements in $S\setminus\langle a\rangle$ are $a^{2i}b$, for $0\leq i\leq 2^{n-2}-1$, of order 2, or $a^{2i+1}b$, for $0\leq i\leq 2^{n-2}-1$, of order 4. Moreover, we have the following useful properties:
  • (i)
    $S$ possesses precisely three maximal subgroups, respectively the cyclic group $C=\left\langle a\right\rangle$, the generalized quaternion group $Q=\left\langle a^{2},ab\right\rangle$, and the dihedral group $D=\left\langle a^{2},b\right\rangle$. More generally, the subgroups of $S$ are cyclic, generalized quaternion, dihedral or Klein four-groups;
  • (ii)
    the center of $S$ is $Z(S)=\left\langle a^{2^{n-2}}\right\rangle$, which has order 2, and $S/Z(S)\cong D_{2^{n-2}}$, the dihedral group of order $2^{n-1}$;
  • (iii)
    the maximal cyclic subgroups of $C$, $D$ and $Q$ all concide with $\Phi(S)=[S,S]=<a^{2}>\cong C_{2^{n-2}}$, and we have $S/\Phi(S)=S/[S,S]\cong C_{2}\times C_{2}$ ;
  • (iv)
    the normal subgroups of $S$ are the three maximal normal subgroups $C$, $D$ and $Q$, together with all the subgroups of $C$. In particular, every proper normal subgroup of $S$ is either maximal or is cyclic and contained in $[S,S]$.
Notice that, by definition, $n\geq 4$, or, in other words, $S$ has order at least $16$. For more details on the properties of a semidihedral group, see, e.g., [1, Lemma 1, p. 9].
3. Proof of Theorem 2.1
In this section, we prove Theorem 2.1, beginning with several lemmas needed to show that any finite group with a semidihedral Sylow 2-subgroup satisfies Property (1).
Throughout this section, we assume that $G$ is a minimal counterexample to Theorem 2.1. In particular, we suppose that $G$ admits a class-preserving, non-inner Coleman automorphism of order a power of 2, which we denote by $\varphi$, without further mention. Since a Sylow 2-subgroup of $G$ is semidihedral, and its subgroups, together with the subgroup itself, are cyclic, dihedral, or generalized quaternion, then the Sylow 2-subgroups of subgroups and quotients of $G$ are of these types. Hence, by the minimality of $G$, [14, Proposition 4.7], and the main theorem of [15], proper quotient, and proper subquotient of $G$ satisfies Property (1).
Let us denote by $S$ a Sylow 2-subgroup of $G$, and let $a,y\in G$ be such that $\left\langle a\right\rangle$ is the maximal cyclic subgroup of $S$ not containing $y)$. In other words, $y$ has order 2 or 4.
Remark 2.
Notice that any proper quotient $G/N$ of $G$ must have even order. Otherwise, $N$ would contain a Sylow 2-subgroup $S=SD_{2^{n}}$, and by Lemma 2.2, since $\varphi$ is assumed to be a non-inner class-preserving Coleman automorphism of order a power of 2, $\varphi$ would induce a class-preserving Coleman automorphism on $N$ of order a power of 2. This contradicts the minimality of $G$. In particular, no proper normal subgroup of $G$ contains $S$.
Let $F=F(G)$ denote the Fitting subgroup of $G$, and let $F^{*}=F^{*}(G)=FE$ be the generalized Fitting subgroup of $G$, where $E=E(G)$ is the layer of $G$. See [9, Chapter 1, Section 4] for more details on the properties of $F$ and $F^{*}$ that we will use in this paper. In particular, we will use that $\prod_{p\in\pi(G)}O_{p}(G), FF^{*}$ is a self centralizing subgroup and is a central product of $F$ and $E$, and $Z(E)$ is an abelian normal subgroup of $G$ contained in $F$.
The proof of Theorem 2.1 relies on several auxiliary lemmas.
Lemma 3.1.
If $G$ is a minimal counterexample to Theorem 2.1, then $O_{2^{\prime}}(F)$ is non-trivial.
Proof.
Suppose, by way of contradiction, that $F$ is a 2-group; that is, $F=O_{2}(G)$. By Remark 1, we can modify $\varphi$ by an inner automorphism so that it fixes $S$ element-wise.
If $F=F^{*}$, then $C_{G}(F)\leq F$. Since $F$ is a 2-group, it follows that $F\leq S$ and so $F$ is fixed element-wise by $\varphi$. Therefore, $\varphi$ induces the identity on $G/F$, and hence, by Lemma 2.5, $\varphi\in\operatorname{Inn}(G)$, which is a contradiction.
Hence, we can assume $E\neq 1$. Suppose $Z(E)=1$, it then follows that $F^{*}=F\times E$. If $E$ contains a cyclic Sylow 2-subgroup, then by the Burnside’s normal $p$-complement theorem [6, Theorem II, Section 243], it has a normal 2-complement, and hence, by Feit-Thompson theorem [8], it is solvable. This is impossible since $E$ is a product of non-abelian simple groups. Therefore, $E$ cannot contain a cyclic Sylow 2-subgroup. Since a Sylow 2-subgroup of $F^{*}=F\times E$ is a subgroup of the semidihedral group $S$, and a Sylow 2-subgroup of $E$ cannot be cyclic, it follows that $O_{2}(G)=F=1$, and $F^{*}=E$ is a non-abelian simple group. In [3, 4, 23], Brauer and Suzuki proved that a generalized quaternion cannot be a Sylow 2-subgroup of a finite simple group. Moreover, by Remark 2, a Sylow 2-subgroup of $E$ cannot be a semidihedral group, since $E$ is a normal subgroup of $G$. Therefore a Sylow 2-subgroup of $E$ can be only a dihedral group. In [10] it was proved that the only finite simple groups with a dihedral Sylow 2-subgroup are $\mathrm{PSL}_{2}(q)$, and $\mathrm{A}_{7}$. Noting that every Coleman automorphism of a group is, up to an inner automorphism, a $p$-central automorphism for any prime $p$ dividing the order of the group, it follows from [16, Theorem 14] that there exists a prime $p\in\pi(E)$ such that all $p$-central automorphisms of $E$ are inner. So, also in this case, we can modify $\varphi$ so that $E=F^{*}$ is fixed element-wise. Consequently, $\varphi$ induces the identity on $G/F^{*}$. Since $Z(F^{*})=Z(E)=1$, it follows that $\varphi=\mathrm{id}$, which is a contradiction. Thus we have $1\neq Z(E)\leq F$ is a 2-group.
Let $P$ be a Sylow 2-subgroup of $E<F^{*}=FE$. As shown above, $P$ cannot be cyclic. Moreover, $P$ cannot be a Klein four-group, since it is a subgroup of a semidihedral group and $F$ is a 2-group, hence it is non-abelian. Consequently, $Z(E)=Z(P)\cong C_{2}$. On the one hand, a Sylow 2-subgroup $P/Z(P)$ of $F^{*}/Z(E)$ is a dihedral group, since $P$ is non-cyclic subgroup of a semidihedral group. On the other hand, since, by Remark 2, a Sylow 2-subgroup of $G$ cannot be a normal subgroup, we conclude that if $F\neq 1$, then $O_{2}(G)=F$ must be either a cyclic group or a dihedral group or a generalized quaternion group. We now show that $F^{*}=E$. Recall that $F^{*}=FE$ with $[F,E]=1$ and $F\cap E=Z(E)$. Since $F=O_{2}(G)$ is a $2$-group, a Sylow $2$-subgroup of $F$ is $F$ itself. Now, a Sylow $2$-subgroup of $F^{*}$ is of the form $F\cdot P$, where $P$ is a Sylow $2$-subgroup of $E$ and $F\cap P=Z(E)\cong C_{2}$, and this must be (isomorphic to) a subgroup of the semidihedral group $S$. Since $[F,E]=1$, we have $[F,P]=1$, and so $F\leq C_{S}(P)$. However, since $P$ is a non-cyclic subgroup of $S$ and $Z(S)\cong C_{2}$, it follows that $C_{S}(P)=Z(S)=Z(P)$. Therefore $F\leq Z(P)=Z(E)$. Since we have already established that $1\neq Z(E)\leq F$, we conclude that $F=Z(E)$. In particular, $F\leq E$, and hence $F^{*}=FE=E$. As above, we can modify $\varphi$ so that it fixes $F^{*}$ element-wise. Moreover, we can suppose that $\varphi$ stabilizes $S$ and that $P\leq S$. Hence, since $\varphi$ is a Coleman automorphism, there exists $s\in S$ such that $\varphi|_{S}=\operatorname{\mathrm{conj}}_{s}|_{S}$ and in particular $[P,s]=1$. If $s\not\in P$, then $\left\langle P,s\right\rangle$ is a non-abelian 2-group with center of order at least 4, which is a contradiction. Otherwise, if $s\in P$, then $s\in Z(P)=Z(E)=Z(F^{*})$. This implies that we can assume $\varphi|_{F^{*}}=\mathrm{id}$ and also $\varphi|_{S}=\mathrm{id}$, since $s\in Z(P)=Z(S)$ and $\varphi|_{S}=\operatorname{\mathrm{conj}}_{s}|_{S}$. By Lemma 2.5, it follows that $\varphi$ is an inner automorphism, which is a contradiction. Therefore, $O_{2^{\prime}}(F)\neq 1$. ∎
Lemma 3.2.
If $G$ is a minimal counterexample to Theorem 2.1, then the Frattini subgroup $\Phi(G)$ of $G$ is a 2-group. In particular, the subgroup $O_{2^{\prime}}(F)$ is a product of abelian minimal normal subgroup of $G$ and it has a complement in $G$.
Proof.
Note that $G$ satisfies the hypotheses of Lemma 2.4, since it is a counterexample of Theorem 2.1. Therefore $\Phi(G)$ is a 2-group. By well-known properties of the Frattini subgroup and the Fitting subgroup (see, e.g. [17, Chapter III, 3.14, 4.4, 4.5]), since $\Phi(G)$ is a 2-group, and so $O_{2^{\prime}}(G)\cap\Phi(G)=1$, the second part of the claim also follows. ∎
Lemma 3.3.
Let $G$ be a minimal counterexample to Theorem 2.1. If $M$ is a minimal normal subgroup of $G$ contained in $O_{2^{\prime}}(F)$, then $M$ has a complement $K$ in $G$, and it is possible to modify $\varphi$ so that $K$ is fixed element-wise by $\varphi$. Moreover, if $N\neq M$ is a normal subgroup of $G$, then there exists a 2-element $k\in K$ such that
  • $k$ inverts the elements of $M$;
  • the image of $k$ in $G/MN$ is a central involution; and
  • $\varphi|_{M}=\operatorname{\mathrm{conj}}_{k}|_{M}$.
Proof.
By hypothesis, $M$ is a minimal normal subgroup contained in $O_{2^{\prime}}(F)$. In particular, since $M$ is normal, it is a union of conjugacy classes, and so $\varphi(M)=M$; moreover, by Lemma 3.2, $O_{2^{\prime}}(F)$ is a product of abelian minimal normal subgroups of $G$, and so, by [12, Proposition 1], $M$ has a complement $K$ such that $K\varphi=K$.
Since $K\cong G/M$, the minimality of $G$ implies that $\varphi$ induces an inner automorphism on $G/M$. Hence, there exists $h\in K$ such that $\mathrm{conj}_{h}^{-1}\varphi$ induces the identity on $G/M$; that is, for every $l\in K$, we have $l(\mathrm{conj}_{h}^{-1}\varphi)=lm$ for some $m\in M$. Since both $\varphi$ and $\mathrm{conj}_{h}$ preserve $K$ (recall that $K\varphi=K$ and $h\in K$), the left-hand side belongs to $K$. As $G=M\rtimes K$ and $M\cap K=1$, it follows that $m=1$. Therefore, $\mathrm{conj}_{h}^{-1}\varphi$ fixes $K$ element-wise. By Remark 1, it is then possible to modify $\varphi$ by an inner automorphism so that $K$ is fixed element-wise by $\varphi$.
Moreover, since $M$ is contained in a Sylow $p$-subgroup $Q$ of $G$ and $\varphi$ is a Coleman automorphism, there exists $g\in G$ such that $\varphi|_{Q}=\mathrm{conj}_{g}|_{Q}$. Writing $g=mh$ with $m\in M$ and $h\in K$, since $M$ is abelian and normal we obtain $\varphi|_{M}=\mathrm{conj}_{h}|_{M}$. Replacing $h$ by its $2$-part (which is possible since $\varphi$ has $2$-power order, and so $\mathrm{conj}_{h}|_{M}$ also has $2$-power order, implying that the odd-order part of $h$ acts trivially on $M$), we get a $2$-element $k\in K$ such that $\varphi|_{M}=\mathrm{conj}_{k}|_{M}$. Note that $k$ is determined by the Coleman property and does not depend on any choice of normal subgroup.
Now, if $N\neq M$ is a normal subgroup of $G$, then by the minimality of $G$, $\varphi$ induces an inner automorphism on $G/N$; that is, $\varphi$ induces $\mathrm{conj}_{\bar{g}}$ on $G/N$ for some $\bar{g}\in G/N$. Since $\varphi$ has $2$-power order, also the induced automorphism has $2$-power order, so we may replace $\bar{g}$ by its $2$-part and assume that $\bar{g}$ is a $2$-element of $G/N$. Now, $G/N=(MN/N)\rtimes(KN/N)$, and $MN/N$ has odd order (since $M$ has odd order). Therefore, all Sylow $2$-subgroups of $G/N$ are contained in conjugates of $KN/N$, and so every $2$-element of $G/N$ is conjugate to an element of $KN/N$. Up to composing with an inner automorphism (which does not change the class modulo $\mathrm{Inn}(G/N)$), we may assume $\bar{g}=k_{N}N$ for some $2$-element $k_{N}\in K$. A priori, $k_{N}$ depends on $N$. However, since $M$ is a minimal normal subgroup and $N\neq M$, we have $N\cap M=1$, and so $\mathrm{conj}_{k_{N}}|_{M}=\varphi|_{M}=\mathrm{conj}_{k}|_{M}$, which gives $k_{N}k^{-1}\in C_{K}(M)$. In particular, $k_{N}\equiv k\pmod{MN}$. Therefore, the image of $k$ in $G/MN$ is independent of the choice of $N$, and $\varphi|_{M}=\mathrm{conj}_{k}|_{M}$.
Let $g=mh\in G$, with $m\in M$ and $h\in K$. Notice that, on the one hand,
$(gMN)\varphi=(MgN)\varphi=Mk^{-1}gk\,N=k^{-1}mhk\,MN,$
while, on the other hand,
$(gMN)\varphi=(mh)\varphi\,MN=k^{-1}mk\,h\,MN.$
Hence we obtain hkMN=khMNhk\,MN=kh\,MN.
Recalling that $MN=NM$, we have
$(MNkg)\varphi=(NMkmh)\varphi=NMkk^{-1}mkh=MNkh,$
and
$(MNgk)\varphi=(NMmhk)\varphi=(NMhk)\varphi=MNhk.$
It follows that the image of $k$ is central in $G/MN$.
Moreover, note that $gMN=NMg=NMh=hMN$. We have shown that $k,G]\leq MN$, and in particular the image $\bar{k}$ of $k$ in $G/MN$ is central. We now show that $\bar{k}$ is an involution, i.e., $k^{2}\in MN$. Since $k\in S$ (up to conjugation, we may assume $k$ belongs to a Sylow $2$-subgroup $S$ of $G$), and $MN\leq O_{2^{\prime}}(F)$ has odd order, we have $S\cap MN=1$. Therefore, $S$ maps isomorphically onto its image $\bar{S}=SMN/MN$ in $G/MN$, which is a Sylow $2$-subgroup of $G/MN$. In particular, $\bar{S}$ is a semidihedral group (isomorphic to $S$), and so $Z(\bar{S})\cong C_{2}$. Since $\bar{k}\in Z(\bar{S})$ (as $k,G]\leq MN$ implies $\bar{k},\bar{S}]=1$), and $\bar{k}$ is a non-trivial $2$-element (otherwise $k\in MN$, which is impossible since $k$ is a $2$-element and $MN$ has odd order, so $k=1$, contradicting the fact that $\varphi$ is non-inner), it follows that $\bar{k}$ generates $Z(\bar{S})$, and in particular $\bar{k}^{2}=1$, that is, $k^{2}\in MN$. Therefore, the image of $k$ in $G/MN$ is a central involution.
Furthermore, $C_{M}(k)=C_{M}(\varphi)\unlhd G$, and so, since $M$ is minimal normal subgroup of $G$, we must have either $C_{M}(\varphi)=M$ or $C_{M}(\varphi)=1$. Since $\varphi$ does not act trivially on $M$, otherwise it is the identity on both $K$ and $M$, i.e. on $G$, we conclude that $C_{M}(\varphi)=1$. Finally, notice that $\varphi|_{M}=conj_{k}|_{M}$ is a fixed-point-free automorphism acting on th elementary abelian group $M$, whose order is coprime to that of $\varphi$. Hence the action of $\varphi|_{M}$ on the finite vector space $M$ is diagonalizable and, since it is fixed-point-free, every eigenvalue needs to be -1. It follows that $k$ acts by inversion on $M$. ∎
Lemma 3.4.
Let $G$ be a minimal counterexample to Theorem 2.1, and let $M$ be a minimal normal subgroup of $G$ contained in $O_{2^{\prime}}(F)$ with complement $K$ in $G$. Suppose that $\varphi$ fixes $K$ element-wise. Assume that $M$ is not a cyclic group, and let $U$ be a maximal subgroup of $M$. Then $\varphi$ induces an inner automorphism on $N_{G}(U)/U$.
Proof.
Note that $M$ is a subgroup of a Sylow $p$-subgroup $Q$ of $G$. Since $\varphi$ is a Coleman automorphism, we have $\varphi|_{Q}=\operatorname{conj}_{g}|_{Q}$ for some $g\in G$. As $M$ is normal in $G$, it follows that for each $m\in M\leq Q$ we have $m\varphi=g^{-1}mg\in M$. Moreover, as shown above, $g$ acts on $M$ by inversion. Therefore, since inversion is an automorphism of the abelian group $M$ and $U$ is a subgroup of $M$, we have $g^{-1}Ug=U$; that is, $g$ normalizes $U$.
We now show that $\varphi$ induces a Coleman automorphism $\psi$ on $N_{G}(U)/U$. First, observe that $\varphi$ stabilizes both $N_{G}(U)$ and $U$: indeed, $\varphi$ fixes $K$ element-wise and acts on $M$ by inversion (which preserves every subgroup of $M$, in particular $U$ and its normalizer in $G$). Hence $\varphi$ induces a well-defined automorphism $\psi$ of $N_{G}(U)/U$.
Let $q$ be a prime and let $\bar{T}$ be a Sylow $q$-subgroup of $N_{G}(U)/U$.
Case $q=p$. Write $\bar{T}=TU/U$, where $T$ is a Sylow $p$-subgroup of $N_{G}(U)$. Then $T$ is contained in a Sylow $p$-subgroup $Q^{\prime}$ of $G$, and $\varphi|_{Q^{\prime}}=\operatorname{conj}_{g^{\prime}}|_{Q^{\prime}}$ for some $g^{\prime}\in G$ which, by the same argument as above, acts on $M$ by inversion and hence normalizes $U$. In particular, $g^{\prime}\in N_{G}(U)$, and for every $t\in T$ we have
$(tU)\psi=t\varphi\,U=g^{\prime-1}tg^{\prime}\,U=(tU)\operatorname{conj}_{g^{\prime}U}.$
Therefore $\psi|_{\bar{T}}$ coincides with the inner automorphism of $N_{G}(U)/U$ given by conjugation by $g^{\prime}U$.
Case $q\neq p$. Since $G=M\rtimes K$ and $q\neq p$, every Sylow $q$-subgroup of $N_{G}(U)$ is conjugate in $N_{G}(U)$ to a subgroup of $N_{K}(U)=K\cap N_{G}(U)$. Because $\varphi$ fixes $K$ element-wise, it fixes $N_{K}(U)$ element-wise, and therefore $\psi$ acts as the identity on every Sylow $q$-subgroup of $N_{G}(U)/U$ that is the image of a subgroup of $N_{K}(U)$. In particular, $\psi|_{\bar{T}}$ is inner (being the identity).
Since $\psi$ restricts to an inner automorphism on every Sylow subgroup of $N_{G}(U)/U$, we conclude that $\psi$ is a Coleman automorphism of $N_{G}(U)/U$.
By Proposition 2.3, the automorphism $\psi$ is also class-preserving, and hence, by the minimality of $G$, it must be inner. ∎
Lemma 3.5.
Let $G$ be a minimal counterexample to Theorem 2.1. Then any proper cyclic quotient of $G$ has order 2.
Proof.
Since $G/N$ is cyclic with Sylow $2$-subgroup isomorphic to $C_{2}$, we have $|G/N|=2m$ for some positive odd integer $m$. Suppose, by way of contradiction, that $|G/N|>2$, i.e., $m>1$. Let $L$ be the unique normal subgroup of $G$ containing $N$ such that $|G/L|=m$. Since $m>1$, $L$ is a proper normal subgroup of $G$. Moreover, since $|G:L|=m$ is odd, $L$ contains a Sylow $2$-subgroup of $G$, and hence the Sylow $2$-subgroups of $L$ are semidihedral. By the minimality of $G$, $L$ satisfies Property 1. Since $2\nmid|G/L|$, Lemma 2.2(ii) implies that $G$ also satisfies Property 1, which contradicts the assumption that $G$ is a counterexample. Therefore, $m=1$ and $|G/N|=2$. ∎
Lemma 3.6.
Let $G$ be a minimal counterexample to Theorem 2.1, and let $E$ be the layer of $G$. Then $O_{2}(G)E$ is non-trivial.
Proof.
Suppose, by way of contradiction, that $O_{2}(G)E=1$. This implies that $F^{*}=F$ is a $2^{\prime}$-group and, by Lemma 3.2, it is a direct product of abelian minimal normal subgroups $M_{i}$ of $G$, with $1\leq i\leq\ell$. Note that $\ell>1$. Indeed, if $F=M_{1}$, since $C_{G}(F)=C_{G}(F^{*})\leq F^{*}=F$, by Lemma 2.4(iii), it follows that $\varphi|_{M_{1}}\neq\gamma|_{M_{1}}$ for each $\gamma\in\operatorname{Inn}(G)$, which contradicts Lemma 3.3. Throughout the remainder of the proof, I will repeatedly make use of the fact – without further explicit mention – that the number $\ell$ of the $M_{i}$ is strictly greater than zero.
Since $O_{2}(G)E=1$, we have $F=O_{2^{\prime}}(F)$, which is a $2^{\prime}$-group. By Lemma 3.2, $F$ is a direct product of abelian minimal normal subgroups $M_{1},\ldots,M_{\ell}$ of $G$, and $F\cap\Phi(G)=1$, so that $F$ has a complement $H$ in $G$, with $G=F\rtimes H$. Since $F\neq 1$ (by Lemma 3.1), $G/F\cong H$ is a proper quotient of $G$, and so, by the minimality of $G$, $\varphi$ induces an inner automorphism on $G/F$. By Remark 1, we can modify $\varphi$ by an inner automorphism so that it induces the identity on $G/F$; equivalently, $\varphi$ fixes $H$ element-wise.
For each index $i$, we now apply Lemma 3.3 to the minimal normal subgroup $M_{i}$, with complement $K_{i}=\hat{M}_{i}\rtimes H$ and taking $N$ to be any normal subgroup of $G$ different from $M_{i}$ (for instance, $\hat{M}_{i}$). Since $\varphi$ already fixes $H$ element-wise and $H\leq K_{i}$, the $2$-element $h_{i}$ provided by Lemma 3.3 can be chosen in $H$. In particular, $h_{i}$ inverts the elements of $M_{i}$, its image in $G/F\cong H$ is a central involution, and $\varphi|_{M_{i}}=\operatorname{conj}_{h_{i}}|_{M_{i}}$.
By Lemma 3.3, applied with $M=M_{i}$ and $N=\hat{M}_{i}$ (so that $MN=F$), the involution $h_{i}$ maps to a central involution in $G/F\cong H$. Since $G=F\rtimes H$ and $F$ has odd order, we may choose the Sylow $2$-subgroup $S$ to lie in $H$. The projection $H\to G/F$ restricts to an isomorphism on $H$, so $h_{i}$ is itself a central involution in $H$. In particular, $h_{i}\in Z(H)\leq Z(S)$. Since $Z(S)$ has order $2$ and contains a unique involution, namely $a^{2^{n-2}}$, it follows that $h_{i}=a^{2^{n-2}}$ for every $i$. In particular, $h_{i}=h_{j}$ for all $1\leq i,j\leq\ell$.
Now we fix an index $i$, and assume that $M_{i}$ is cyclic, generated by $m_{i}$. Let $K=C_{H}(M_{i})$. Note that $H\leq N_{G}(M_{i})$ implies $C_{H}(M_{i})\unlhd H$. Moreover, since $M_{i}$ is a cyclic minimal normal subgroup of odd order, it follows that $M_{i}\cong C_{p}$ for some prime $p$. Hence, $H/K$ is isomorphic to a subgroup of $C_{p-1}$, and so it is cyclic. Furthermore, $h_{i}$ does not belong to $K$, since it acts by inversion on $M_{i}$, and so, by Lemma 3.5, $H/K\cong C_{2}$, generated by $h_{i}\in H$. Thus, we conclude that $H=K\rtimes\langle h_{i}\rangle$. Since $h_{i}$ maps to a central involution in $H/K$, we have $H=K\times\langle h_{i}\rangle$. Recall that, since $F$ has odd order, a Sylow $2$-subgroup $S$ of $G$ is contained in $H$ (up to conjugacy). The projection $H\to\langle h_{i}\rangle\cong C_{2}$ restricts to a homomorphism $S\to C_{2}$, whose kernel $S\cap K$ has index at most $2$ in $S$. Since $S\cap K$ is a subgroup of $S$ of index at most $2$, it must contain one of the three maximal subgroups of $S$ (the cyclic group $C$, the dihedral group $D$, or the generalized quaternion group $Q$). In all cases, $h_{i}=a^{2^{n-2}}\in[S,S]\leq S\cap K\leq K$. But $h_{i}\in\langle h_{i}\rangle$ and $K\cap\langle h_{i}\rangle=1$, which forces $h_{i}=1$, a contradiction. Therefore, $M_{i}$ cannot be cyclic.
Now we modify $\varphi$ so that it fixes $\hat{M}_{1}H$ element-wise and $\varphi|_{M_{1}}=\operatorname{\mathrm{conj}}_{h_{1}}|_{M_{1}}$, and we choose a maximal subgroup $U$ of $M_{1}$. By Lemma 3.4, $\varphi$ induces an inner automorphism on $N_{G}(U)/U$. Then there exists a non-trivial 2-element $h\in N_{H}(U)$ such that $h,\hat{M}_{1}]=1$ and $h$ acts on $M_{1}/U$ by inversion. Substituting $S$ with a suitable conjugate, if necessary, we may suppose $h\in S$. The same reasoning applied to another minimal normal subgroup, for instance $M_{2}$, shows that there exists a non-trivial 2-element $k\in H$ such that $k,\hat{M}_{2}]=1$, so in particular $k,M_{1}]=1$, and we may suppose that $k\in S$. Indeed, $k$ is a $2$-element, so it is contained in some Sylow $2$-subgroup $S^{\prime}$ of $G$. Since all Sylow $2$-subgroups are conjugate, there exists $g\in G$ such that $S^{\prime}=g^{-1}Sg$. Setting $k^{\prime}=gkg^{-1}\in S$, we have $k^{\prime},M_{1}]=g[k,M_{1}]g^{-1}=1$, since $k,M_{1}]=1$ and $M_{1}$ is normal in $G$. Therefore, we may assume $k\in S$. As seen above, the element $h_{1}$ in $H$, acting on $M_{1}$ by inversion, is a central involution in $S$, and we may assume that it belongs to the maximal cyclic subgroup $C=\langle a\rangle$ of $S=SD_{2^{n}}$. Since $h_{1}$ acts on $M_{1}$ by inversion and $M_{1}$ has no elements of order 2, it follows that $h_{1}$ acts fixed-point free on $M_{1}$. In particular, $h_{1}=a^{2^{n-2}}$. Moreover, $a$ also acts fixed-point free on $M_{1}$ and $k,M_{1}]=1$, and so we may suppose that $k\not\in\left\langle a\right\rangle$, that is, $k=a^{s}b$, for some $s\in\mathbb{N}$ and $b$ as introduced in Notation 2.6. Also, $h=a^{t}b$, for some $t\in\mathbb{N}$, since $h,\hat{M}_{1}]=1$ and $h_{1}$ – which equals $h_{i}$ for all $i$ – acts fixed-point free on $\hat{M}_{1}$. Then $hk^{-1}=a^{t-s}$. Since $h=a^{t-s}k$, and $k$ fixes $M_{1}$ element-wise, it follows that $h$ acts on $M_{1}$ by conjugation with $a^{t-s}$. Since $h$ acts on $M_{1}/U$ by inversion, it follows that $a^{t-s}=h_{1}$. Therefore, $h=a^{t-s}k=h_{1}k$, and this implies that $\varphi|_{F}=\operatorname{\mathrm{conj}}_{h}|_{F}$, which contradicts Lemma 2.4(iii). ∎
Lemma 3.7.
Let $G$ be an minimal counterexample to Theorem 2.1. Then $O_{2}(G)$ is non-trivial.
Proof.
Suppose, by way of contradiction, that $O_{2}(G)=1$. This implies that the order of $F$ is odd, and by Lemma 3.6, $E\neq 1$. Moreover, $Z(E)$ also has odd order. However, since $Z(E)=\Phi(E)\leq\Phi(G)$, and $\Phi(G)$ is a $2$-group, it follows that $Z(E)=1$. Since $Z(E)=1$, the layer $E$ is a direct product of non-abelian simple groups $L_{1},\ldots,L_{r}$. Suppose $r\geq 2$. Then a Sylow $2$-subgroup of $E$ is of the form $T_{1}\times\cdots\times T_{r}$, where each $T_{i}$ is a Sylow $2$-subgroup of $L_{i}$. By Burnside’s normal $p$-complement theorem, a non-abelian simple group cannot have a cyclic Sylow $2$-subgroup, and hence $|T_{i}|\geq 4$ for every $i$. In particular, $T_{1}\times T_{2}$ is an abelian subgroup of $S$ of order at least $16$ and rank at least $2$. However, the only abelian subgroup of $S=SD_{2^{n}}$ of rank $2$ is the Klein four-group, which has order $4$. This is a contradiction. Therefore $r=1$, and $E$ is a non-abelian simple group. By Lemma 3.2, there exists an abelian minimal normal subgroup of $O_{2^{\prime}}(F)$ which has a complement $K$ in $G$. We claim that $E\leq K$. Since $M$ and $E$ are both normal in $G$ and $M\cap E\leq F\cap E=Z(E)=1$, it follows that $[M,E]=1$. Now, consider the projection $\pi\colon G=M\rtimes K\to K$ (with kernel $M$). Since $E\cap M=1$, the restriction $\pi|_{E}$ is injective. Let us define $\sigma\colon E\to M$ such that $e\sigma=e\cdot(e\pi)^{-1}\in M$ for each $e\in E$. Notice that every $e\in E$ can be written as $e=e\sigma\cdot e\pi$. A direct computation shows that the map $\sigma\colon E\to M$ is a crossed homomorphism with respect to the conjugation action of $E\pi$ on $M$. Since $[E,M]=1$, the subgroup $E\pi$ centralizes $M$, and so $\sigma$ is in fact a group homomorphism. As $E$ is perfect (being the layer of $G$) and $M$ is abelian, we have $E\sigma\leq[M,M]=1$. Therefore $\sigma=1$, which means $e=e\pi\in K$ for all $e\in E$, and hence $E\leq K$. Moreover, we can suppose that $\varphi$ fixes $K$ element-wise and acts on $M$ by inversion.
Suppose that $M$ is not cyclic. Let $U$ be a maximal subgroup of $M$. By Lemma 3.4, $\varphi$ induces an inner automorphism on $N_{G}(U)/U$. Hence, there exists a 2-element $g\in K$ such that $(h\varphi)h^{-g}\in U$ for each $h\in N_{G}(U)$. In particular, since $E\leq K$ and $\varphi$ fixes $K$ element-wise, we have $[E,g]=1$, and $g$ acts on $M/U$ by inversion. Let $P$ be a Sylow 2-subgroup of $E$. As shown in the proof of Lemma 3.1, $P$ is not cyclic, $[P,g]=1$, and $\langle P,g\rangle$ is a 2-group with center of order greater than or equal to 4, which is impossible. Therefore, $M$ is cyclic, and, by Lemma 3.5, $K/C_{K}(M)\cong C_{2}$.
Since $K/C_{K}(M)\cong C_{2}$ (as shown above by Lemma 3.5), and $\bar{k}\notin C_{K}(M)/E$ (as $k$ acts on $M$ by inversion), we have $K/E=\langle\bar{k}\rangle\cdot(C_{K}(M)/E)$ with $\langle\bar{k}\rangle\cap(C_{K}(M)/E)=1$. Moreover, by Lemma 3.3, $\bar{k}$ is a central involution in $G/ME\cong K/E$, and so the product is direct: $K/E=\langle\bar{k}\rangle\times(C_{K}(M)/E)$.
Assume that $C_{K}(M)=E$, that is, $G/ME=\left\langle\bar{k}\right\rangle\cong C_{2}$. Let $1\neq m\in M$ and $e\in E$. Then there exists $c\in K$ such that $(me)\varphi=(me)^{c}$, that is, $m^{-1}=m\varphi=c^{-1}mc$ and $e=e\varphi=c^{-1}ec$. It follows that $c=hk$, for some $h\in E$, and so $\varphi|_{E}=\operatorname{\mathrm{conj}}_{k}|_{E}$ is a class-preserving automorphism of $E$. Hence, $\operatorname{\mathrm{conj}}_{k}|_{E}\in\operatorname{Inn}(E)$, considering the structure of $E$ described in the proof of Lemma 3.1 and recalling that a class-preserving automorphism, up to an inner automorphism, is $p$-central for every prime $p$ dividing the order of $E$. In other words, there exists a 2-element $e\in E$ such that $\operatorname{\mathrm{conj}}_{k}|_{E}=\operatorname{\mathrm{conj}}_{e}|_{E}$, and, in particular, $[k,e]=1$. This implies that $ke^{-1}$ is a 2-element not belonging to $E$, and it centralizes a Sylow 2-subgroup of $E$. Note that this leads to a contradiction: indeed, $E$ must be either a dihedral or a generalized quaternion group, and the subgroup generated by $ke^{-1}$ centralizes a Sylow 2-subgroup of $E$ while having trivial intersection with $E$. However, this is not possible, since the Sylow 2-subgroups of $G$ are semidihedral. Therefore, we conclude that $C_{K}(M)/E\neq 1$.
Let $P$ be a Sylow 2-subgroup of $C_{K}(M)/E$. A Sylow $2$-subgroup of $K/E$ has the form $\langle\bar{k}\rangle\times P\cong C_{2}\times P$, and is a quotient of the semidihedral group $S$ (since $F$ has odd order, a Sylow $2$-subgroup of $G$ maps onto a Sylow $2$-subgroup of $G/E\cong M\rtimes(K/E)$, and the $2$-part concentrates in $K/E$). The only quotient of $S=SD_{2^{n}}$ that admits a direct factor isomorphic to $C_{2}$ is $S/[S,S]\cong C_{2}\times C_{2}$. Indeed, any proper quotient of $S$ is either cyclic or dihedral or isomorphic to $C_{2}\times C_{2}$, and among these, only $C_{2}\times C_{2}$ contains a direct factor $C_{2}$ as a proper subgroup. It follows that the Sylow $2$-subgroup of $K/E$ is $C_{2}\times C_{2}$, and in particular $P\cong C_{2}$. Hence, a Sylow 2-subgroup of $G/E=MK/E$ is an elementary abelian 2-group of order 4, and so it is $S/[S,S]$. It follows that a Sylow 2-subgroup of $E$ is $[S,S]$, which is cyclic. However, $E$ cannot contain a cyclic Sylow 2-subgroup, and so this leads to a contradiction. Therefore, we conclude that $O_{2}(G)\neq 1$. ∎
Lemma 3.8.
Let $G$ be a minimal counterexample to Theorem 2.1, and let $S=SD_{2^{n}}$ be a semidihedral Sylow $2$-subgroup of $G$. Then $Z(S)$ is normal in $G$. In particular, $Z(S)=O_{2}(G)$
Proof.
First note that, by Lemma 3.7 and Remark 2, $O_{2}(G)$ is a non-trivial proper normal subgroup of $S$.
Suppose that $O_{2}(G)$ is the maximal dihedral subgroup $D$ of $S$ or the maximal generalized quaternion subgroup $Q$ of $S$. Note that $Z(S)=Z(D)=Z(Q)$, and since the center is characteristic in both $D$ and $Q$, it follows that $Z(S)$ is normal in $G$ in either case.
Now suppose that $O_{2}(G)$ is one of the cyclic subgroups contained in the maximal cyclic subgroup $C$ of $S$. If $O_{2}(G)=Z(S)$, we are done. Otherwise, in each of the other cases – recalling that every subgroup of a cyclic group is characteristic – it follows again that $Z(S)$ is normal in $G$.
Finally, suppose $O_{2}(G)\neq Z(S)$. It follows that $O_{2}(G)$ has order at least 4.
Let $M$ be a minimal normal subgroup of $G$ contained in $O_{2^{\prime}}(F)$, whose existence is guaranteed by Lemma 3.1, and let $K$ be a complement of $M$ in $G$, as provided by Lemma 3.3. We may assume that $\varphi$ fixes $K$ elementwise, and there exists a $2$-element $\tilde{k}\in K\cap S$ such that $\varphi|_{M}=\operatorname{conj}_{\tilde{k}}|_{M}$ and $\tilde{k},G]\leq MZ(S)$.
Now let $z\in O_{2}(G)\setminus Z(S)$ and $1\neq m\in M$. Since $\varphi$ fixes $K$ elementwise and $\varphi|_{M}=\operatorname{conj}_{\tilde{k}}|_{M}$, we have $(mz)\varphi=m^{\tilde{k}}\cdot z$. On the other hand, $(mz)\varphi$ must be conjugate to $mz$ (as $\varphi$ is class-preserving), so there exists $g\in G$ such that $(mz)\varphi=g^{-1}(mz)g$. Writing $g=m^{\prime}k$ with $m^{\prime}\in M$ and $k\in K$, we note that $M,O_{2}(G)]=1$ (since $M\leq O_{2^{\prime}}(F)$ and $O_{2}(G)$ are both normal in $G$ with coprime orders, so $M,O_{2}(G)]\leq M\cap O_{2}(G)=1$). Together with the fact that $M$ is abelian, this gives $(mz)^{m^{\prime}}=mz$, and hence $(mz)^{g}=(mz)^{k}$, so we may take $g=k\in K$. Since $\varphi$ has $2$-power order and acts on $M$ via $\tilde{k}$, we may further assume that $g$ is a $2$-element of $K$.
Then there exists a 2-element $g\in K$ such that $(mz)\varphi=(mz)^{g}$, that is, $m^{-1}=m\varphi=m^{g}$ and $z=z\varphi=z^{g}$. It follows that $\left\langle O_{2}(G),g\right\rangle$ is a non-cyclic abelian 2-group of order at least 8. We now prove the latter claim.
First, we show that $g$ centralizes $O_{2}(G)$. From the equality $z=z^{g}$, we know that $g$ centralizes $z$. Since we are in the case where $O_{2}(G)$ is a cyclic subgroup of the maximal cyclic subgroup $C=\langle a\rangle$ of $S$ (the cases $O_{2}(G)\in\{D,Q\}$ having already been treated at the beginning of this proof), and $O_{2}(G)$ has order at least $4$, the element $z\in O_{2}(G)\setminus Z(S)$ generates a subgroup $\langle z\rangle$ of index at most $2$ in $O_{2}(G)$. Now, $g$ is a $2$-element of $S$ that centralizes $\langle z\rangle$, and hence it induces an automorphism of $O_{2}(G)$ that fixes a subgroup of index at most $2$ pointwise. Since $O_{2}(G)$ is cyclic of $2$-power order, the only automorphism of $O_{2}(G)$ that fixes a subgroup of index at most $2$ elementwise is the identity. Therefore $g$ centralizes $O_{2}(G)$, and so $\langle O_{2}(G),g\rangle$ is abelian.
Next, we show that $\langle O_{2}(G),g\rangle$ is non-cyclic. Suppose, for a contradiction, that $\langle O_{2}(G),g\rangle$ is cyclic. Then $g\in O_{2}(G)$, since a cyclic $2$-group has a unique subgroup of each order. But $O_{2}(G)\leq C\trianglelefteq S$ and $M,O_{2}(G)]=1$, so $g\in O_{2}(G)$ would imply $m^{g}=m$ for every $m\in M$, contradicting $m^{g}=m^{-1}\neq m$. Therefore $g\notin O_{2}(G)$, and $\langle O_{2}(G),g\rangle$ is a non-cyclic abelian $2$-group. Since $|O_{2}(G)|\geq 4$ and $g\notin O_{2}(G)$, the group $\langle O_{2}(G),g\rangle$ has order at least $8$.
Hence $\langle O_{2}(G),g\rangle$ is a non-cyclic abelian $2$-group of order at least $8$, which is impossible since all abelian subgroups of the semidihedral group $S$ have order at most $2^{n-1}$ and are cyclic or isomorphic to $C_{2}\times C_{2}$. Hence, $O_{2}(G)=Z(S)$. ∎
Before stating the following lemma, we recall some previously established properties of a minimal counterexample to Theorem 2.1. By Lemma 3.1 and Lemma 3.3, there exists a minimal normal subgroup of $G$ contained in $O_{2^{\prime}}(F)$, which has a complement $K$ in $G$. Modifying $\varphi$ according to Lemma 3.3, we may assume that $\varphi$ fixes $K$ element-wise. Furthermore, by Lemma 3.7 and again Lemma 3.3, there exists a 2-element $\tilde{k}\in K$ – which may be taken to lie in $S\leq K$ – such that $\tilde{k},G]\leq MO_{2}(G)$, and both $\varphi$ and $\tilde{k}$ act on $M$ by inversion. Note that, by Lemma 3.4 and Lemma 3.8, we also conclude that $\tilde{k},G]\leq MZ(S)$; in particular, $\tilde{k},S]\leq Z(S)$.
Lemma 3.9.
Let $G$ be a minimal counterexample to Theorem 2.1, and let $M$ be a minimal normal subgroup of $G$ contained in $O_{2^{\prime}}(F)$. Then $M$ is cyclic.
Proof.
The structure of the semidihedral group $S=SD_{2^{n}}$ implies that if $1\neq[k,S]\leq Z(S)$, then $\tilde{k}\in\left\langle a\right\rangle$ and has order 4; that is, $\tilde{k}=a^{2^{n-3}}$. Let $s=y$ be an element of order 2 or 4 in $S\setminus C$, where $C$ is the maximal normal cyclic subgroup of $S$, and let $P=\langle\tilde{k},s\rangle$. Note that $Z(P)=Z(S)$.
Assume, by way of contradiction, that $M$ is not cyclic. Note that $s^{2}=1$ or $s^{2}=a^{2^{n-2}}\in Z(S)$; hence $s^{2},M]=1$, since, by Lemma 3.8, $Z(S)\cong C_{2}$ is normal in $G$. In particular, there exists a maximal subgroup $U$ of $M$ such that $P$ normalizes $U$, that is, $P\leq N_{K}(U)$. By Lemma 3.4, $\varphi$ induces an inner automorphism on $N_{K}(U)M/U$. Thus, there exists a $2$-element $g\in N_{K}(U)$ such that $(h\varphi)h^{-g}\in U$ for all $h\in N_{K}(U)M$. Since $\varphi$ fixes $K$ element-wise and $P\leq N_{K}(U)$, we have $p\varphi=p$ for every $p\in P$. By Lemma 3.4, $(p\varphi)p^{-g}\in U$ for every $p\in P$, that is, $pp^{-g}\in U\leq M$. On the other hand, since both $p$ and $g$ belong to $K$, we have $pp^{-g}\in K$. Therefore $pp^{-g}\in M\cap K=1$, which gives $[g,p]=1$ for every $p\in P$. Hence $[g,P]=1$, and so $\langle P,g\rangle$ is a $2$-group.
If $g\in P$, then $g\in Z(P)=Z(S)$, and so $g$ would act trivially on $M$; however, this is impossible since, by Lemma 3.4, $g$ acts on $M$ by inversion on $M/U$. If $g\notin P$, then $\left\langle P,g\right\rangle$ would be a non-abelian 2-group with center of order at least 4, which is a contradiction, since $Z(P)=Z(S)\cong C_{2}$. Therefore, we conclude that $M$ is cyclic. ∎
Now we are ready to conclude the proof of Theorem 2.1.
Proof of Theorem 2.1.
Let us consider the notations introduced above, and let $G$ be a minimal counterexample to Theorem 2.1. By Lemma 3.9, a minimal normal subgroup $M$ of $G$ contained in $O_{2^{\prime}}(F)$ is cyclic; in particular, $M\cong C_{p}$ for some prime $p$. It follows that $K/C_{K}(M)$ is cyclic, and thus, by Lemma 3.5, we have $K/C_{K}(M)\cong C_{2}$.
By the beginning of the proof of Lemma 3.9, we know that $\tilde{k}\in K$ belongs to the cyclic subgroup of order 4 in $S$, which is normal in $S$ and lies in $S,S]\cong C_{2^{n-2}}$. This leads to a contradiction, since $K/C_{K}(M)\cong C_{2}$, and so its Sylow 2-subgroups must be isomorphic to either $S/C_{2^{n-1}}$, $S/D_{2^{n-2}}$, or $S/Q_{2^{n-1}}$. In all these cases, $S,S]\leq C_{K}(M)$, implying that $K$ acts trivially on $M$. However, we know that $\tilde{k}\in K$ acts on $M$ by inversion. Hence, we obtain a contradiction.
This completes the proof of the main theorem. ∎
Acknowledgements
The author acknowledges the funding support from MUR-Italy via PRIN 2022RFAZCJ "Algebraic methods in Cryptanalysis". The author wishes to express sincere gratitude to the referee for the careful reading of the manuscript and for the valuable suggestions, which have significantly contributed to improving the clarity of the exposition.